OpenStudy (anonymous):

write this equation in slope intercept form: 8=12y+6x

OpenStudy (anonymous):

8=12y+6x -12y -12y -12y+8=6x -8 -8 -12y=6x-8 /-12 /-12 y=-1/2+2/3

OpenStudy (anonymous):

ooops, forgot..... TADA

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thanks a lot bro

OpenStudy (anonymous):

no problem(:

OpenStudy (anonymous):

can you answer another problem i have?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

6(x+3y) = 36

OpenStudy (anonymous):

k, imma draw this one out

OpenStudy (anonymous):

go ahead :)

OpenStudy (anonymous):